OpenStudy (anonymous):

Find the present value of the deposit. In order to have $9000 at 10% compounded monthly for 4 years.. Need help with this.. at least the formula. Thanx!! ;)

OpenStudy (campbell_st):

the formula is \[A = P(1 + r)^n\] A = future value, P = Principal (amount invested) r = interest rate as a decimal n = number of time periods.. is the interest rate 10% per annum or 10% per month...?

OpenStudy (kropot72):

\[9000=P(1+\frac{0.1}{12})^{(4\times 12)}=P(1.008333333)^{48}\] Now solve to find P

OpenStudy (anonymous):

why I got this answer $13,404.19? and the right is $6,042.89. What I am doing wrong? Thanks!

OpenStudy (anonymous):

@campbell_st is per month..

OpenStudy (kropot72):

@rocal2 If you solve the equation that I posted you will get the correct answer. Do you have a problem with solving the equation?

OpenStudy (anonymous):

Yes, I guess I am pluging something wrong.. What I am doing is 9000(1+0.1/12)^48. It is this is correct?

OpenStudy (campbell_st):

well looking at the correct answer @kropot72 method is correct...

OpenStudy (kropot72):

\[P=\frac{9000}{(1+\frac{0.1}{12})^{48}}\]

OpenStudy (anonymous):

Yes I am agree with that.. I am pretty sure I am doing something wrong..

OpenStudy (anonymous):

@kropot72 Got it now... thank you!!

OpenStudy (kropot72):

You're welcome :)